Assume that Betty's current dividend Do is $1.00; it is expected to grow 15% the first year, 20% the second year, 10% the third year, and return to its long-run constant growth rate of 4%. Cost of equity, or required rate of return is 7%.

  • What is the stock’s horizon terminal value?
  • What is the stock's intrinsic stock price today?
Answer #1

1.

Stock's horizon terminal value=1*1.15*1.2*1.1*1.04/(7%-4%)=$52.62

2.

Stock's intrinsic price today=1*1.15/1.07+1*1.15*1.2/1.07^2+1*1.15*1.2*1.1/1.07^3+1*1.15*1.2*1.1/1.07^3*1.04/(7%-4%)=$46.48